### v1

# An Improved Two-Step Attack on CRYSTALS-Kyber

Abstract.
After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber has successfully been selected and drafted for standardization from the mid of 2022. It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase. In this paper, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time. In the first step, we use the correlation power analysis (CPA) attack to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces. The CPA attack is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy. In the second step, we adopt the lattice attack to recover s based on the results of CPA. The success rate is largely built up by constructing a trail-and-error method. We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine. Additionally, in that case, we only cost at most $60$ CPA guess values for a group and $15$ power traces for a guess.
Lattice-based cryptography, CRYSTALS-Kyber, Side-channel attack, Power analysis
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 2024 ACM/IEEE International Conference on Computer-Aided Design; October 27–31, 2024; NEW JERSEY, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Traditional public-key cryptography Rivest-Shamir-Adleman (RSA) algorithm (rivest1978method, ) and elliptic-curve cryptography (ECC) (koblitz1987elliptic, ) rely on the computational intractability of integer decomposition and discrete logarithm problems respectively. However, concerns have been raised with the emergence of quantum computing because they can be cracked in polynomial time by Shor’s algorithm (shor1994algorithms, ), and thereby revealing the security of existing cryptographic algorithms is insufficient. Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process with the aim of standardizing quantum-resistant cryptographic algorithms in 2016 (avanzi2017crystals, ). By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and a key encapsulation mechanism (KEM) algorithm in the third round (alagic2022status, ). CRYSTALS-Kyber (bos2018crystals, ) is that only KEM.
Kyber is a lattice-based cryptographic algorithm constructed based on the module-learning with errors (M-LWE) problem. Even in quantum computing, the M-LWE problem is considered to be secure (avanzi2019crystals, ). It should be noted that the mathematical security of Kyber has been widely recognized by the cryptography community. However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security.
Kocher et al. first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ). Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs such as simple power analysis (SPA) attack, template attack (TA), and correlation power analysis (CPA) attack (avanzi2019crystals, ). Different kinds of SCAs for Kyber have gradually emerged. In terms of SPA, Xu et al. successfully conducted such attack on the inverse number theoretic transform (INTT) of the reference implementation of Kyber512 and the pqm4 implementation by constructing specific ciphertext pairs, recovering a coefficient of a secret key costs $8\sim 960$ power traces (xu2021magnifying, ). For TA, Works in (chari2003template, ; choudary2014efficient, ; mu2022voltage, ) analyzed power traces collected from a large number of devices using belief propagation techniques to construct attack templates, enabling the recovery of keys from individual power traces. All of them require tremendous extra data to establish good templates. For CPA, most of previous works (ravi2020generic, ; ueno2022curse, ; shen2023find, ) are chosen-ciphertext attacks (CCAs). The latest work in (shen2023find, ) can significantly reduce the power traces compared to prior works by using an efficient two-step scheme to deal with the imperfect SCA oracles. However, since the NIST PQC KEMs are CCA secure, all of them need extra efforts such as plaintext checking to verify the final results. In (yang2023chosen, ), Yang et al. tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces within a few minutes. Meanwhile, they have demonstrated that the chosen ciphertext CPA attack can improve the efficiency in some degree. Nevertheless, They adopted the original CPA method and the rest of coefficients of the secret key are needed to be recovered separately in the same way. In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu recently presented a novel two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack. They constructed a lower dimension of M-LWE problem in the NTT process based on the guess values of CPA and directly calculated the full key. Two hundreds simulated traces are used in their experiments to recover the secret key with about 20 minutes on a 16-core machine for Kyber512. However, they only conducted computer simulations. In this paper, we take the solution of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
The proposed two steps can be overviewed as Fig. 1. In this first step, we apply an enhanced CPA attack to recover parts of coefficients of secret key by exploiting the combined correlation between the modified Hamming weight (HW) for some intermediate values and the power consumption of the decryption process in Kyber, specifically the point-wise multiplication covering a secret polynomial and a ciphertext. In this way, some of the secret coefficients can be recovered using a small amount of power traces in number theoretic transform (NTT) domain. In the second step, We take the points of the lattice attack used in (kuo2023lattice, ) and construct a trail-and-error algorithm to recover the entire secret key. Our main contributions can be summarized as follows:
  • We use both Pearson and Kendall’s rank correlation coefficients (Kendall’s tau) and modify the leakage model to improve the accuracy of CPA attacks.
  • Based on lattice attack, we construct a trail-and-error algorithm to improve the success rate.
  • We combine the two steps of attacks together and apply to Kyber512 on ARM Cortex-M4. Experimental results show that the proposed method only costs $38\sim 60$ CPA guess values and about $9$ minutes on a $16$-core machine to recover the 128 coefficients of a secret key s, much faster than the state-of-the-art.
The rest of this paper is structured as follows. Section 2 provides preliminary content. Section 3 presents the traditional attacks and our attack method. Experimental results are described in Section 4. Section 5 summarizes the paper.
Figure 1. The flowchart of the proposed attack.
2. Preliminaries
In this section, we will introduce the notations, the principles of the LWE/M-LWE problem, NTT in Kyber, CPA Attack, and Kendall’s tau.
2.1. Notations
The ring of integers modulo the prime number $q$ is denoted as $\mathbb{Z}_{q}$. Each polynomial ring $\mathcal{R}_{q}$ = $\mathbb{Z}_{q}[x]/(x^{n}+1)$ with moduli $x^{n}+1$ and $q$ has $n$ coefficients. The Greek symbols $\beta_{\eta}$ and $\mathcal{U}$ denote the centered binomial distribution with parameter $\eta$ and the uniform distribution, respectively. The vectors are represented by bold lowercase letters, such as $\mathbf{a}$, and the vectors in the NTT domain have a hat added to them, such as $\hat{\mathbf{a}}$. The matrices are represented by bold uppercase letters, such as $\mathbf{A}$.
Table 1. Parameter in CRYSTALS-Kyber
$k$ $n$ $q$ $d_{u}$ $d_{v}$ $\eta_{1}$ $\eta_{2}$
Kyber512 2 256 3329 10 4 3 2
Kyber768 3 256 3329 10 4 2 2
Kyber1024 4 256 3329 11 5 2 2
2.2. LWE and M-LWE
Regev et al. introduced the LWE problem (regev2009lattices, ), which forms the basis for several NIST PQC candidates. As show in (1), the LWE problem involves recovering the invariant secret vector $\mathbf{s}$ from $m$ equations, where fixed $\mathbf{s}\leftarrow\beta_{\eta}(\mathbb{Z}_{q}^{n})$.
(1) $(\mathbf{a}_{i},b_{i}=\mathbf{a}_{i}^{\top}\mathbf{s}+e_{i})\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q},$
where $\mathbf{a}_{i}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and the error vector $e_{i}\leftarrow\beta_{\eta}(\mathbb{Z}_{q})$. The essence of the M-LWE problem lies in replacing the ring $\mathbb{Z}_{q}$ in the above LWE problem with the polynomial ring $\mathcal{R}_{q}$, and the error distribution is $\beta_{\eta}(\mathcal{R}_{q})$. Thus, an M-LWE sample $(\mathbf{a}_{i},b_{i})\leftarrow\mathcal{U}(\mathcal{R}_{q}^{k\times 1}\times\mathcal{R}_{q})$ can be represented as:
(2) $(\mathbf{a}_{i},b_{i}=\mathbf{a}_{i}^{\top}\mathbf{s}+e_{i})\in\mathcal{R}_{q}^{k\times 1}\times\mathcal{R}_{q},$
where $\mathbf{a}_{i}\leftarrow\mathcal{U}(\mathcal{R}_{q}^{k\times 1})$ and the error vector $e_{i}\leftarrow\beta_{\eta}(\mathcal{R}_{q})$. Thus, a set of m M-LWE samples can be integrated as:
(3) $(\mathbf{A},\mathbf{b}=\mathbf{A}\mathbf{s}+\mathbf{e})\in\mathcal{R}_{q}^{m\times k}\times\mathcal{R}_{q}^{m}.$
2.3. NTT in CRYSTALS-Kyber
Kyber is the only standardized KEM of PQC algorithms established by NIST in the third round (moody2021nist, ) and is built on the M-LWE problem. It offers three NIST security levels: Kyber512 corresponding to Level 1, Kyber768 to Level 3, and Kyber1024 to Level 5. The specific parameters are shown in Table 1.
The public-key encryption (PKE) scheme involved in KEM of Kyber consists of three stages: key generation, encryption, and decryption. In the key generation stage, the public key $pk$ is constructed as $\mathbf{A}\mathbf{s}+\mathbf{e}$, where $\mathbf{A}$ is the sampling polynomial matrix. $\mathbf{s}$ and $\mathbf{e}$ are the secret key and noise, respectively, both of which are polynomial vectors sampled from the $\beta_{\eta_{1}}$. During the encryption stage, the message $m$ is encrypted into ciphertext $c$, where $c$ is formed by compressing a polynomial vector $u$ and an array $v$ and concatenating them. In the decryption stage, the receiver extracts $u$ and $v$ from the ciphertext $c$, and then utilizes $\mathbf{s}$ to perform corresponding operations to recovery the message $m$. Algorithm 1 illustrates the decryption process of Kyber. The KEM protocol is an extension of the PKE protocol with re-encryption. The application of the Fujisaki-Okamoto transformation (fujisaki1999enhance, ) to an IND-CPA-secure PKE results in an IND-CCA2-secure KEM.
In Kyber, polynomial multiplication is a fundamental operation that is frequently used in the encryption and decryption processes. By performing multiplication calculations on the polynomial converted to the NTT domain, the polynomial multiplication on NTT can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus accelerating the speed of the entire encryption and decryption process. Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512_{\text{th}}$. Therefore, the modulus $x^{n}+1$ in Kyber can only be partially factored into $n/2$ quadratic polynomials, with odd and even coefficients calculated respectively. It is also considered that NTT is a linear transformation, the specific formula is described as follows:
[…]
where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}_{n\times n}$ is a reduced integer matrix. Similarly, INTT can be represented in the same way.
Algorithm 1 Kyber.CPAPKE.Dec(sk,c):decryption
1: Input: Secret key $sk\in\mathcal{B}^{12\cdot k\cdot n/8}$
2: Input: Ciphertext $c=(\mathbf{u},v)\in\mathcal{B}^{d_{u}\cdot k\cdot n/8+d_{v}\cdot n/8}$
3: Output: Message $m\in\mathcal{B}^{32}$
4:$\mathbf{u}:=\text{Decompress}_{q}(\text{Decode}_{d_{u}}(c),d_{u})$
5:$v:=\text{Decompress}_{q}(\text{Decode}_{d_{v}}(c+d_{u}\cdot k\cdot n/8),d_{v})$
[…]
8:return $m$
2.4. Correlation Power Analysis Attack
SCAs on cryptographic devices can be broadly categorized into invasive, non-invasive, and semi-invasive attacks. Non-invasive attacks, such as the CPA attack, can compromise secret keys without disrupting the operation of the cryptographic devices. In the following, we will delve into the principles and steps involved in CPA attack in summary.
In practical SCAs, physical phenomena such as the power consumption and the electromagnetic radiation are often observed (verbauwhede2010secure, ). The CPA attack is a widely used method in SCAs, which relies on the correlation between power models such as HW or Hamming distance (HD) and actual power consumption (joy2011side, ). The higher the correlation guess, the greater the likelihood that the key with higher correlation guess is the correct one. A traditional CPA attack can generally be divided into four steps.
Firstly, select an intermediate value $\mathit{f}(d,k)$ computed by the device’s encryption algorithm as the attack point, where $d$ is a known partial ciphertext or message and $k$ is a partial key. Secondly, measure the actual power consumption of the device. When the device encrypts or decrypts $D$ different messages or ciphertexts, we record the actual power consumption $\mathbf{T}_{D\times T}$, where $\mathbf{T}_{ij}$ denotes the power consumption of the $i^{th}$ plaintext or ciphertext at time $j$. Thirdly, calculate the hypothetical intermediate values and map them to the real power consumption. The hypothesis intermediate $\mathbf{V}=\left[f(d_{i},gk_{j})\right]_{D\times K}$ is computed for all guess keys $gk$ and mapped to the power consumption $\mathbf{H}_{D\times K}$. Finally, calculate the Pearson correlation coefficient (PCC) between $\mathbf{T}$ and $\mathbf{H}$ to obtain the correlation coefficient matrix $\mathbf{P}_{K\times T}$. According to the largest value in $\left|\mathbf{P}\right|$, we determine the correct part of the private key $k_{c}$ and time. The full secret key can be recovered by repeating the above procedure for each partial secret key.
2.5. Kendall’s Rank Correlation Coefficient
The Kendall’s tau is a non-parametric measure of the association between two random variables. It quantifies the degree of concordance in the rankings of two variables, i.e., whether they are consistently ranked in terms of their values (abdi2007kendall, ). Kendall’s tau $\tau$ ranges from $-1$ to $1$, where $\tau=1$ indicates a perfect direct association, $\tau=-1$ indicates a perfect disassociation, and $\tau=0$ indicates no agreement in rankings between the two variables.
[…]
where $c$ and $d$ denote the numbers of concordant and discordant pairs, while $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
Listing 1: Base multiplication in Kyber.
1void basemul(int16_t r[2],
2 int16_t zeta,
3 const int16_t s[2],
4 const int16_t u[2])
5{
6 r[0] = fqmul(s[1], u[1]);
7 r[0] = fqmul(r[0], zeta);
8 r[0] += fqmul(s[0], u[0]);
9 r[1] = fqmul(s[0], u[1]);
10 r[1] += fqmul(s[1], u[0]);
11}
3. Proposed Improved Two-Step Attack
It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We will detail the basis of the attack analysis and our proposed three improvements in the following.
3.1. Basis of Attack Analysis
The goal of this paper is to explore the physical security of Kyber KEM under random-ciphertext CPA attack. The purpose of the attack is to obtain a long-term secret key. We assume that an attacker has access to a device which is running Kyber decryption and can enter arbitrary ciphertext into the device. In addition, they can also capture the electromagnetic radiation of the device to obtain the power traces.
In subsection 2.3, we have introduced NTT used in Kyber briefly, which only possesses $n_{\text{th}}$ roots of unity because the modulus polynomial $(x^{n}+1)$ can only be factored into $n/2$ linear polynomials. Consequently, multiplication involves the multiplication of linear polynomials rather than simple point-wise multiplication in the NTT domain. Due to the properties of incomplete NTT in Kyber, the full secret key of Kyber can be divided into $2k(k=2,3,4)$ groups.
The test vector leakage assessment (TVLA) methodology can be used to find side-channel leakage points and the t-test method is one of the commonly used methods in TVLA (wu2022efficient, ). Several t-test results from previous works have indicated the potential leakage points during the decryption process in Kyber (yang2023chosen, ; cryptoeprint:2022/058, ). One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$. Additionally, this operation takes place in the quotient ring $\mathbb{Z}_{q}[x]/(x^{2}-\xi)$. In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512. Moreover, each calculation of $basemul$ is independent from the others.
From Listing 1, we can observe that there are two steps related to the partial secret key coefficient $s_{0}$ and three steps related to the partial secret key coefficient $s_{1}$. The inputs of a $basemul$ are $s_{0}$, $s_{1}$, $u_{0}$, and $u_{1}$, and outputs are $r_{0}$ and $r_{1}$. Five multiplications and two additions are involved in the listing. Obviously, the result is related to two secret key coefficients. All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber. In a practical CPA attack, suppose that the attacker aims to recover the coefficient $s_{0}$, $r_{0}$ can be selected as the intermediate value. Similarly, if the attacker aims to recover the coefficient $s_{1}$, $r_{1}$ can be selected as the intermediate value. In addition, in order to recover the full secret key coefficients of a group, the attacker would need to perform 128 CPA attacks. Therefore, $128\times 2k$ CPA attacks are required for the full secret key.
Listing 2: Modified HW in the CPA Attack.
1def HW(r0):
2 if r0 < 0:
3 r0 = calculate_complement(r0, 16)
4 H_W = r0.count("1")
5 else:
6 H_W = bin(r0).count("1")
7 return H_W
8
9def calculate_complement(r0, 16):
3.2. Modified Leakage Model for CPA
Modeling the power leakage is the basis for a CPA attack, and its accuracy directly determines the success rate of the attack. HW model counts the number of “1” in intermediate values, which is one of the most representative linear power models. Similarly, HD model records the amount of ”1$\rightarrow$”0” or ”0$\rightarrow$”1” transitions as power leakage. The difference of HW and HD model is given as:
(6) $HD\left(r_{0},s_{1}\right)=HW\left(r_{0}\oplus s_{1}\right).$
Typically, when simulating the energy consumption of a microprocessor, the HW model is commonly employed (zhao2023side, ), as is the case in this particular study.
The arithmetic logic units (ALUs) and multiplier units (MU) in ARM processors are often optimized for two’s complement operations. Most arithmetic and logical instructions in the processor’s instruction set are designed based on two’s complement calculations. Therefore, we fine-tune the HW power leakage model according to this property to directly avoid the complementary false positive. We compute the HW of the complement of the intermediate instead of directly computing the HW of the intermediate, as shown in Listing 2, where the function calculate_complement computes the $16bit$ complement. The modified power leakage model can map to the hypothetical power consumption values more accurately, thus reducing the number of power traces required for CPA attacks. As show in Fig. 2, correct secret key coefficient in the original HW leakage model is drowned in all the guessed key coefficients, while the coefficient can be picked out in the modified model effectively by using a small number of 15 power traces.
Figure 2. When $D=15$ during a $basemul$ call, PCC values for all guessed keys of the original HW model (top) and the modified HW model (bottom).
3.3. Optional Kendall’s tau for CPA
The goal of this improvement is to find the correct key coefficients using 15 or even fewer power traces while ensuring the efficiency and success rate of the attack. In scenarios where the number of power traces is limited, the calculated PCC value across all guessed keys tends to be exceptionally high. Consequently, it becomes challenging to accurately recover even a portion of the secret key coefficients using solely the traditional CPA attack, and attempting to reconstruct 128 or the entire set of secret key coefficients becomes even more daunting.
Our solution is as follows. After computing the PCC correlation matrix for each $basemul$, we set the threshold to 0.9 or even higher, adjusting it based on the values of $\mathbf{P}$. For each $basemul$, if there are more than $f$ correlations exceeding the threshold, we consider them as candidate coefficients $\mathbf{k_{h}}=(k_{h_{0}},...,k_{h_{f-1})}$. We then map the hypothetical intermediate values of the $f$ candidate coefficients to candidate hypothetical power consumption value matrix $\mathbf{H^{\prime}}_{D\times f}$. Next, we calculate the Kendall’s tau between the hypothetical power value matrix $\mathbf{H^{\prime}}$ and the actual power matrix $\mathbf{T}$. For each column of $\mathbf{H^{\prime}}$, we compute the Kendall’s tau with each column of the actual power matrix. Finally, we obtain a $f\times T$ Kendall coefficient matrix $\mathbf{K_{d}}$, from which we select the candidate key with the highest correlation coefficient as the correct key coefficient for that $basemul$. On the other hand, we skip the cases that have higher PCC correlations and smaller candidates.
PCC is used in the traditional CPA attack to calculate only the linear correlation between the actual power consumption and the hypothetical power consumption, which is not sufficient. With only a few power traces, it is possible to directly select some of the correct key coefficients even if only Kendall’s tau is calculated, but the calculation time of Kendall’s tau is 12×12$\times$12 × longer than that of PCC. Therefore we combine Kendall’s tau with PCC. After calculating PCC, the guessed keys coefficient exceeding a certain threshold is selected as the coefficients of candidate keys. Then, the Kendall’s tau values of these coefficients of candidate keys are calculated so that the coefficient with the higher value is accepted. In this way, false positives can be quickly eliminated.
3.4. Trail-and-Error Lattice Attack after CPA
As mentioned earlier, each group of the secret key consists of 128 secret key coefficients including 128 even-index coefficients or 128 odd-index coefficients. As introduced in (kuo2023lattice, ), suppose that an attacker wants to exploit the side-channel leakage point $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$ and has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered.
Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients. INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients. Similarly, $\mathbf{N}_{B}=[\mathbf{n}_{b_{0}},\ldots,\mathbf{n}_{{b}_{127-sr}}]$, and $\hat{\mathbf{s}}_{iB}=[\hat{\mathbf{s}}_{b_{0}},\ldots,\hat{\mathbf{s}}_{{b}_{127-sr}}]^{\top}$.
In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known. Let $\mathbf{t}=\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$, $\mathbf{A}=-\mathbf{N}_{B}$, and $\mathbf{s}^{\prime}=\hat{\mathbf{s}}_{iB}$. Then, we obtain $\mathbf{t}=\mathbf{A}\cdot\mathbf{s}^{\prime}+\mathbf{s}_{i}\mod q$. This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller. We firstly view it as a bounded distance decoding (BDD)/unique shortest vector problem (uSVP) lattice problem, and specific algorithm to solve the updated LWE problem is given by Algorithm 2 (kannan1987minkowski, ), where the matrixs $\mathbf{BDD}$ and $\mathbf{B}_{kan}$ are summarized as follows:
(7) $\mathbf{BDD}=\left[\begin{array}[]{cc}\mathbf{I}_{sr}&\mathbf{A}^{\prime}\\ \mathbf{0}&q\mathbf{I}_{n-sr}\\ \end{array}\right],$
[…]
$\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose. This process yields the shortest vector containing the error vector $\mathbf{s}_{i}^{\top}$, denoted as $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$. Finally, We obtain the actual secret key with a length of 128 as shown in Step 3 of Algorithm 2.
But it is worth noting that the norm of vector $\mathbf{w}$ is $\sqrt{\|s_{i}\|^{2}+1}\approx\sqrt{n}\sigma_{s}$, which must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the uSVP problem to be solved. For Kyber512, when $sr\geq 39$, it is possible to correctly recover the length of 128 odd/even index coefficients through lattice attack. For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered. In other words, the lattice attack can tolerate at most $89/90$ recovered incorrect secret key coefficients for Kyber512 and Kyber768/1024 respectively, where $38/39$ of the coefficient are randomly selected from $128$ recovery coefficients.
We elaborate on the theoretical principles of the lattice attack, which provides a fault-tolerant for recovering the full set of 128 keys. Leveraging this method, we propose a more flexible and versatile approach for its application. The proposed trail-and-error lattice attack is shown in Listing 3, where $n$ represents the number of CPA attacks, and $\mathbb{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trails, respectively.
Listing 3: Trail-and-Error Lattice Attack.
1index = [1, 2, ..., 37, 38]
2other_index = [39, 40, ..., n-2, n-1]
3guess_sk = [sk_bu0, ..., sk_bu38]
4other_sk = [sk_bu39, ..., sk_bu(n-1)]
5ken = [k_bu0, ..., k_bu38]
6other_ken = [k_bu39, ..., k_bu(n-1)]
7
8result = Lattice_attack(index, guess_sk)
9result = (M @ result)
10if guess_sk == [result[j] for j in index]:
11 return result
12else:
13 for i in count:
14 updata_index(index)
15 updata_sk(guess_sk)
16 result = Lattice_attack(index, guess_sk)
17 result = (M @ result)
18 result_target = [result[j] for j in index]
19 if guess_sk == result_target:
20 return result
21 break
22
23def Lattice_attack(a, b):
24def updata_index():
25def updata_sk():
Algorithm 2 Kannan’s embedding technique.
1: Input: An LWE instance $\mathbf{t}=\mathbf{A}\mathbf{s}+\mathbf{s}_{i}\bmod q$
2: Output: the short error vector $\mathbf{s}_{i}\in\mathcal{R}_{q}^{128}$
3:Construct the lattice $\Lambda(\text{BDD})$ generated by matrix $\mathbf{BDD}$;
4:Reduce BDD to the SVP problem by rescaling $\mathbf{BDD}$ to the basis matrix $\mathbf{B}_{kan}$;
5:Use lattice algorithm (LLL or BKZ) to derive the short vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ from $\mathbf{B}_{kan}$;
6:return $\mathbf{s}_{i}$
One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA attacks instead of $128$ CPA attacks. Certainly, the more CPA attacks are carried out, the higher the success rate of this method will be, but its efficiency will also slow down. The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trail-and-error method. It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trail and lead to the overall recovery inefficient.
For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively. Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
Figure 3. The experimental equipment.
4. Experiments
4.1. Experimental Setup
Our attacks are equally successful for different security levels of Kyber because they use the same $basemul$, whose algorithm is provided by Listing 1. For simplicity, we only take Kyber512 into consideration in the following. We run the reference implementation of Kyber512 from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz. We capture the power traces using a Pico 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe. The sampling rate is set to $500$ MSa/s. The experimental equipment is shown in Fig. 3.
According to the introduction in Section 2.4 that the selection of the intermediate value of CPA attack should be related to the secret key, combined with the analysis mentioned in Section 3.1, we choose $fqmul(k,u_{1})$ as the intermediate value of our attack, and all the intermediate values in the following experiments are the same values.
Figure 4. When $D=15$, the PCC results during a certain $basemul$ call.
4.2. Experiment Results
Step1: CPA Attack
As aforementioned, the CPA attack is divided into two stages: the capturing stage and the modeling computation stage. Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
For the capturing stage, as introduced in Section 3.1, we carry out the CPA attack at the NTT operation which calls $basemul$ functions 128 times. In our CPA experiments, we input random ciphertexts and capture tremendous electromagnetic radiations during those calls.
To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments. The first one is shown in Fig. 2 of Section 3.2 and we can see that the new model can reinforce the PCC value of the correct key. The second one shown in the following is to demonstrate its superiority in identifying the complementary false positive.
Figure 5. When $D=15$, the PCC results (a) and Ken results (b) in a certain $basemul$. When $D=11$, the PCC results (c) and Ken results (d) in the same $basemul$.
When the number of random ciphertexts $D$ is set to 15, the relationship of the maximum absolute values $\left|PCC\right|$ to the guess keys is shown in the above of Fig. 4. We can directly pick out the correct key 2280 and its complementary false positive 1049 (equal to $3329-2280$). This is a general phenomenon. As mentioned in (kuo2023lattice, ), the attacker would run into some problems because the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ both would have high $PCC$ values since $HW(s_{0})$ and $HW(q-s_{0})$ are highly correlated. So, we have modified the leakage model as described in Section 3.2 to escape such false positive. The relationship of the PCC values to the trace points is shown in the bottom of of Fig. 4, where there are 3329 curves and the yellow and red curves present whose guess keys are 2280 and 1049, respectively. It can be seen that the maximum absolute value of PCC of the complementary false positive has a negative sign. We can only focus on the positive PCCs to avoid such false positive.
We briefly explain the reason of this phenomenon in the following. Let $s_{0}$ be the correct one. Then $s_{0}^{\prime}=3329-s_{0}$ represents the false positive. The $fqmul$ function in $basemul$ is an operation that multiplies two numbers and then is reduced by $3329$ in Kyber. The intermediate values for the two guessed key coefficients are summarized as follows:
(9) $r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$
(10) $r_{1}^{\prime}=\text{fqmul}(s_{0}^{\prime},u_{1})$ $=((3329-s_{0})\times u_{1})\,mod\,q$
$=-(s_{0}\times u_{1})\,mod\,q.$
Hypothetical power consumption values $MHW(r_{1})$ and $MHW(r_{1}^{\prime})$ exhibit a negative correlation, while correct $MHW(r_{1})$ will naturally exhibit a positive correlation. Obviously, we can figure them out by taking their signs into consideration.
To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process. Figure 5 (a) shows the points of the maximum absolute PCC results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates $f$ after PCC is equal to 4. It can be seen that the correlations of four coefficients $564$, 1687, 2010, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to 0.930 and their correlation values are very close. Very close correlations can lead to an inability to distinguish the correct coefficient. This is an another kind of false positives. If the attacker just picks one of these four at random, the accuracy is only about 25%, which is still regarded as a failure. Therefore, we use the method mentioned in Section 3.3 to compensate for such case. First of all, we compute the Kendall coefficient matrix of these four candidate coefficients. It should be noted that the time for this computation is negligible when comparing to the complete PCC computing. Figure 5 (b) shows the points of the four maximum absolute values $\left|Kendall’s\,tau\right|$ to the numbers of power traces. It shows that Kendall’s tau expands the differences between correct coefficient and false positives. As the number of power traces increases, the value of correct coefficient tends to be more stable while those of false positives decrease quickly. We can easily distinguish the correct key 564 when the number of power traces equals 15 and output it as the final result. In another word, an attacker can take the correct coefficient 564 for this $basemul$ call using only 15 power traces. Similarly, we reduce the number of power traces to 11 step by step and implement the PCC and Kendall’s tau computations. The corresponding results are shown in Fig. 5 (c) and (d). It can be seen that the correct key 564 can almost not be recognized. We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections. Note that if we continue reducing the number of power traces, we cannot figure it out any more.
Step2: Trail-and-Error Lattice Attack
We conduct the trail-and-error lattice attack on a sixteen-thread server. The parameter of the $block$ size in the BKZ reduction is set to 50 and the value of $max\_loops$ to 8. We collect the CPAs of the first sixty $basemuls$ calls.
Figure 6 shows the relationships of the success rate and time to the number of power traces representing in the blue curve and groups of histograms, respectively. The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times. It can be seen that when the number of power traces is no less than 15, all the experimental attacks are successful. Those cases with traces number smaller than 15 can be compensated to some degree by running more trails of the lattice attack. Meanwhile, according to the decomposed time histograms in Fig. 6, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA attack is almost unchanged although the traces number is doubled. The main reason could be the different utilization ratios of the multi-threading server. Note that the CPA computations are independently executed with the different power traces while the trail computations of the lattice attack are executed in serial.
Therefore, we can summarize that when the success rate is large enough (close to 1), we can only cost 15 power traces for a guess/call and about 9 minutes to recover the full key on Kyber512. So, the total number of required power traces is $15\times 60=900$, much smaller than that of the state-of-the-art random ciphertext CPA (yang2023chosen, ) which requires $20\times 128=2560$ power traces. If we slightly increasing the traces number of the guess, the required time can be reduced close to 5 minutes, much faster than the recorded 20 minutes reported in (kuo2023lattice, ).
5. Conclusion
In this paper, we propose an efficient two-step attack including an enhanced CPA attack and a trail-and-error lattice attack for Kyber. In the CPA step, we modify the power leakage model to be more suitable for ARM Cortex-M4 architecture and further filter the candidate keys from the PCC results by using the Kendall’s rank correlation coefficient. The accuracy of finding the correct key is significantly improved. In the lattice step, we construct a trail-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time. Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads. Because the $basemul$ function called by different security levels of Kyber is the same, our work can be directly applied to the other parameters. Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.
Figure 6. The success rate and efficiency of our attack in twenty experiments.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grant 62104097, in part by the Key Research Plan of Jiangsu Province of China under Grant BE2022098, and in part by the Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001.

### v2

@@
-An Improved Two-Step Attack on CRYSTALS-Kyber
+An Improved Two-Step Attack on Kyber
 Abstract.
-After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber has successfully been selected and drafted for standardization from the mid of 2022.
+After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024.
 It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase.
-In this paper, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time.
-In the first step, we use the correlation power analysis (CPA) attack to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces.
-The CPA attack is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy.
+In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time.
+In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces.
+The CPA is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy.
 In the second step, we adopt the lattice attack to recover s based on the results of CPA.
-The success rate is largely built up by constructing a trail-and-error method.
+The success rate is largely built up by constructing a trial-and-error method.
 We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine.
@@
 1. Introduction
-Traditional public-key cryptography Rivest-Shamir-Adleman (RSA) algorithm (rivest1978method, ) and elliptic-curve cryptography (ECC) (koblitz1987elliptic, ) rely on the computational intractability of integer decomposition and discrete logarithm problems respectively.
-However, concerns have been raised with the emergence of quantum computing because they can be cracked in polynomial time by Shor’s algorithm (shor1994algorithms, ), and thereby revealing the security of existing cryptographic algorithms is insufficient.
-Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process with the aim of standardizing quantum-resistant cryptographic algorithms in 2016 (avanzi2017crystals, ).
-By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and a key encapsulation mechanism (KEM) algorithm in the third round (alagic2022status, ).
-CRYSTALS-Kyber (bos2018crystals, ) is that only KEM.
-Kyber is a lattice-based cryptographic algorithm constructed based on the module-learning with errors (M-LWE) problem.
-Even in quantum computing, the M-LWE problem is considered to be secure (avanzi2019crystals, ).
-It should be noted that the mathematical security of Kyber has been widely recognized by the cryptography community.
-However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security.
+WITH the fast development of quantum computing, modern public-key cryptographic algorithms are encountering significant threats.
+Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms.
+By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation (alagic2022status, ).
+CRYSTALS-Kyber is that only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024 (Announcing2024, ).
+It should be marked that the mathematical security of Kyber has been widely recognized by the cryptography community.
+However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security (nagata2021physical, ; zhao2023side, ; ravi2024side, ).
 Kocher et al.
 first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ).
-Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs such as simple power analysis (SPA) attack, template attack (TA), and correlation power analysis (CPA) attack (avanzi2019crystals, ).
-Different kinds of SCAs for Kyber have gradually emerged.
-In terms of SPA, Xu et al.
-successfully conducted such attack on the inverse number theoretic transform (INTT) of the reference implementation of Kyber512 and the pqm4 implementation by constructing specific ciphertext pairs, recovering a coefficient of a secret key costs $8\sim 960$ power traces (xu2021magnifying, ).
-For TA, Works in (chari2003template, ; choudary2014efficient, ; mu2022voltage, ) analyzed power traces collected from a large number of devices using belief propagation techniques to construct attack templates, enabling the recovery of keys from individual power traces.
-All of them require tremendous extra data to establish good templates.
-For CPA, most of previous works (ravi2020generic, ; ueno2022curse, ; shen2023find, ) are chosen-ciphertext attacks (CCAs).
-The latest work in (shen2023find, ) can significantly reduce the power traces compared to prior works by using an efficient two-step scheme to deal with the imperfect SCA oracles.
-However, since the NIST PQC KEMs are CCA secure, all of them need extra efforts such as plaintext checking to verify the final results.
-In (yang2023chosen, ), Yang et al.
-tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces within a few minutes.
-Meanwhile, they have demonstrated that the chosen ciphertext CPA attack can improve the efficiency in some degree.
-Nevertheless, They adopted the original CPA method and the rest of coefficients of the secret key are needed to be recovered separately in the same way.
-In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu recently presented a novel two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack.
-They constructed a lower dimension of M-LWE problem in the NTT process based on the guess values of CPA and directly calculated the full key.
-Two hundreds simulated traces are used in their experiments to recover the secret key with about 20 minutes on a 16-core machine for Kyber512.
-However, they only conducted computer simulations.
-In this paper, we take the solution of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
-The proposed two steps can be overviewed as Fig.
+Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs, especially to the correlation power analysis (CPA).
+Recent years, several CPA attacks on Kyber (ravi2020generic, ; ueno2022curse, ; shen2023find, ; yang2023chosen, ; kuo2023lattice, ) have been proposed by researchers.
+In 2020, Ravi et al (ravi2020generic, ).
+used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time.
+Then, Ueno et al (ueno2022curse, ).
+proposed the side channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key.
+The work in (shen2023find, ) deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber512 at the cost of about 1619 power traces, but there are still 40% error coefficients.
+Contrast to previous works, in (yang2023chosen, ), Yang et al.
+tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces.
+Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber512.
+So, they at least need 5120 power traces for a key.
+In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack.
+They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of M-LWE problem.
+The correctness was verified by using computer simulations.
+In this paper, we take the idea of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
+Figure 1. The flowchart of the two-step attack.
+The proposed two-step attack is summarized in Fig.
 1.
-In this first step, we apply an enhanced CPA attack to recover parts of coefficients of secret key by exploiting the combined correlation between the modified Hamming weight (HW) for some intermediate values and the power consumption of the decryption process in Kyber, specifically the point-wise multiplication covering a secret polynomial and a ciphertext.
-In this way, some of the secret coefficients can be recovered using a small amount of power traces in number theoretic transform (NTT) domain.
-In the second step, We take the points of the lattice attack used in (kuo2023lattice, ) and construct a trail-and-error algorithm to recover the entire secret key.
-Our main contributions can be summarized as follows:
-  • We use both Pearson and Kendall’s rank correlation coefficients (Kendall’s tau) and modify the leakage model to improve the accuracy of CPA attacks.
-  • Based on lattice attack, we construct a trail-and-error algorithm to improve the success rate.
-  • We combine the two steps of attacks together and apply to Kyber512 on ARM Cortex-M4.
-Experimental results show that the proposed method only costs $38\sim 60$ CPA guess values and about $9$ minutes on a $16$-core machine to recover the 128 coefficients of a secret key s, much faster than the state-of-the-art.
-The rest of this paper is structured as follows.
-Section 2 provides preliminary content.
-Section 3 presents the traditional attacks and our attack method.
-Experimental results are described in Section 4.
-Section 5 summarizes the paper.
-Figure 1. The flowchart of the proposed attack.
+In the first step, we modify the leakage model to better align our CPA with the arithmetic logic unit.
+Additionally, we introduce an optional correlation coefficient, the Kendall rank correlation, to address the nonlinearity and robustness issues of the original CPA in real-world attacks.
+In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key, and optimize the parameters of the algorithm.
+By employing this two-step attack, we achieve significant improvements in both efficiency and accuracy in practice.
 2. Preliminaries
-In this section, we will introduce the notations, the principles of the LWE/M-LWE problem, NTT in Kyber, CPA Attack, and Kendall’s tau.
-2.1. Notations
-The ring of integers modulo the prime number $q$ is denoted as $\mathbb{Z}_{q}$.
-Each polynomial ring $\mathcal{R}_{q}$ = $\mathbb{Z}_{q}[x]/(x^{n}+1)$ with moduli $x^{n}+1$ and $q$ has $n$ coefficients.
-The Greek symbols $\beta_{\eta}$ and $\mathcal{U}$ denote the centered binomial distribution with parameter $\eta$ and the uniform distribution, respectively.
-The vectors are represented by bold lowercase letters, such as $\mathbf{a}$, and the vectors in the NTT domain have a hat added to them, such as $\hat{\mathbf{a}}$.
-The matrices are represented by bold uppercase letters, such as $\mathbf{A}$.
-Table 1. Parameter in CRYSTALS-Kyber
-$k$ $n$ $q$ $d_{u}$ $d_{v}$ $\eta_{1}$ $\eta_{2}$
-Kyber512 2 256 3329 10 4 3 2
-Kyber768 3 256 3329 10 4 2 2
-Kyber1024 4 256 3329 11 5 2 2
-2.2. LWE and M-LWE
-Regev et al.
-introduced the LWE problem (regev2009lattices, ), which forms the basis for several NIST PQC candidates.
-As show in (1), the LWE problem involves recovering the invariant secret vector $\mathbf{s}$ from $m$ equations, where fixed $\mathbf{s}\leftarrow\beta_{\eta}(\mathbb{Z}_{q}^{n})$.
-(1) $(\mathbf{a}_{i},b_{i}=\mathbf{a}_{i}^{\top}\mathbf{s}+e_{i})\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q},$
-where $\mathbf{a}_{i}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and the error vector $e_{i}\leftarrow\beta_{\eta}(\mathbb{Z}_{q})$.
-The essence of the M-LWE problem lies in replacing the ring $\mathbb{Z}_{q}$ in the above LWE problem with the polynomial ring $\mathcal{R}_{q}$, and the error distribution is $\beta_{\eta}(\mathcal{R}_{q})$.
-Thus, an M-LWE sample $(\mathbf{a}_{i},b_{i})\leftarrow\mathcal{U}(\mathcal{R}_{q}^{k\times 1}\times\mathcal{R}_{q})$ can be represented as:
-(2) $(\mathbf{a}_{i},b_{i}=\mathbf{a}_{i}^{\top}\mathbf{s}+e_{i})\in\mathcal{R}_{q}^{k\times 1}\times\mathcal{R}_{q},$
-where $\mathbf{a}_{i}\leftarrow\mathcal{U}(\mathcal{R}_{q}^{k\times 1})$ and the error vector $e_{i}\leftarrow\beta_{\eta}(\mathcal{R}_{q})$. Thus, a set of m M-LWE samples can be integrated as:
-(3) $(\mathbf{A},\mathbf{b}=\mathbf{A}\mathbf{s}+\mathbf{e})\in\mathcal{R}_{q}^{m\times k}\times\mathcal{R}_{q}^{m}.$
 2.1. NTT in CRYSTALS-Kyber
-Kyber is the only standardized KEM of PQC algorithms established by NIST in the third round (moody2021nist, ) and is built on the M-LWE problem.
-It offers three NIST security levels: Kyber512 corresponding to Level 1, Kyber768 to Level 3, and Kyber1024 to Level 5.
-The specific parameters are shown in Table 1.
-The public-key encryption (PKE) scheme involved in KEM of Kyber consists of three stages: key generation, encryption, and decryption.
-In the key generation stage, the public key $pk$ is constructed as $\mathbf{A}\mathbf{s}+\mathbf{e}$, where $\mathbf{A}$ is the sampling polynomial matrix.
-$\mathbf{s}$ and $\mathbf{e}$ are the secret key and noise, respectively, both of which are polynomial vectors sampled from the $\beta_{\eta_{1}}$.
-During the encryption stage, the message $m$ is encrypted into ciphertext $c$, where $c$ is formed by compressing a polynomial vector $u$ and an array $v$ and concatenating them.
-In the decryption stage, the receiver extracts $u$ and $v$ from the ciphertext $c$, and then utilizes $\mathbf{s}$ to perform corresponding operations to recovery the message $m$.
-Algorithm 1 illustrates the decryption process of Kyber.
-The KEM protocol is an extension of the PKE protocol with re-encryption.
-The application of the Fujisaki-Okamoto transformation (fujisaki1999enhance, ) to an IND-CPA-secure PKE results in an IND-CCA2-secure KEM.
-In Kyber, polynomial multiplication is a fundamental operation that is frequently used in the encryption and decryption processes.
-By performing multiplication calculations on the polynomial converted to the NTT domain, the polynomial multiplication on NTT can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus accelerating the speed of the entire encryption and decryption process.
-Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512_{\text{th}}$.
-Therefore, the modulus $x^{n}+1$ in Kyber can only be partially factored into $n/2$ quadratic polynomials, with odd and even coefficients calculated respectively.
-It is also considered that NTT is a linear transformation, the specific formula is described as follows:
-[…]
-where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}_{n\times n}$ is a reduced integer matrix. Similarly, INTT can be represented in the same way.
-Algorithm 1 Kyber.CPAPKE.Dec(sk,c):decryption
-1: Input: Secret key $sk\in\mathcal{B}^{12\cdot k\cdot n/8}$
-2: Input: Ciphertext $c=(\mathbf{u},v)\in\mathcal{B}^{d_{u}\cdot k\cdot n/8+d_{v}\cdot n/8}$
-3: Output: Message $m\in\mathcal{B}^{32}$
-4:$\mathbf{u}:=\text{Decompress}_{q}(\text{Decode}_{d_{u}}(c),d_{u})$
-5:$v:=\text{Decompress}_{q}(\text{Decode}_{d_{v}}(c+d_{u}\cdot k\cdot n/8),d_{v})$
-[…]
-[…]
-8:return $m$
-2.4. Correlation Power Analysis Attack
-SCAs on cryptographic devices can be broadly categorized into invasive, non-invasive, and semi-invasive attacks.
-Non-invasive attacks, such as the CPA attack, can compromise secret keys without disrupting the operation of the cryptographic devices.
-In the following, we will delve into the principles and steps involved in CPA attack in summary.
-In practical SCAs, physical phenomena such as the power consumption and the electromagnetic radiation are often observed (verbauwhede2010secure, ).
-The CPA attack is a widely used method in SCAs, which relies on the correlation between power models such as HW or Hamming distance (HD) and actual power consumption (joy2011side, ).
-The higher the correlation guess, the greater the likelihood that the key with higher correlation guess is the correct one.
-A traditional CPA attack can generally be divided into four steps.
-Firstly, select an intermediate value $\mathit{f}(d,k)$ computed by the device’s encryption algorithm as the attack point, where $d$ is a known partial ciphertext or message and $k$ is a partial key.
-Secondly, measure the actual power consumption of the device.
-When the device encrypts or decrypts $D$ different messages or ciphertexts, we record the actual power consumption $\mathbf{T}_{D\times T}$, where $\mathbf{T}_{ij}$ denotes the power consumption of the $i^{th}$ plaintext or ciphertext at time $j$.
-Thirdly, calculate the hypothetical intermediate values and map them to the real power consumption.
-The hypothesis intermediate $\mathbf{V}=\left[f(d_{i},gk_{j})\right]_{D\times K}$ is computed for all guess keys $gk$ and mapped to the power consumption $\mathbf{H}_{D\times K}$.
-Finally, calculate the Pearson correlation coefficient (PCC) between $\mathbf{T}$ and $\mathbf{H}$ to obtain the correlation coefficient matrix $\mathbf{P}_{K\times T}$.
-According to the largest value in $\left|\mathbf{P}\right|$, we determine the correct part of the private key $k_{c}$ and time.
-The full secret key can be recovered by repeating the above procedure for each partial secret key.
-2.5. Kendall’s Rank Correlation Coefficient
-The Kendall’s tau is a non-parametric measure of the association between two random variables.
-It quantifies the degree of concordance in the rankings of two variables, i.e., whether they are consistently ranked in terms of their values (abdi2007kendall, ).
-Kendall’s tau $\tau$ ranges from $-1$ to $1$, where $\tau=1$ indicates a perfect direct association, $\tau=-1$ indicates a perfect disassociation, and $\tau=0$ indicates no agreement in rankings between the two variables.
-[…]
-where $c$ and $d$ denote the numbers of concordant and discordant pairs, while $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
+Computing polynomial multiplication in NTT domain can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus significantly improving the speed of the entire encryption and decryption process in Kyber.
+Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512^{\text{th}}$.
+Therefore, 256 coefficients are divided into two groups, namely the odd and even groups, for NTT calculations.
+The even coefficients in NTT domian are described as follows:
+(1) $\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$
+where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}$ is a $128\times 128$ matrix with reduced integers.
+Similarly, INTT can be represented in the same way and the inverse matrix is denoted as $N$.
 Listing 1: Base multiplication in Kyber.
@@
 11}
+2.2. Correlation-Based Attack Techniques
+The conventional CPA attack is a non-invasive side-channel attack that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys.
+It identifies the correct key by finding the highest correlation between measured and hypothetical values.
+Pearson correlation coefficient (PCC) is efficient and well-suited for detecting linear relationships between the hypothetical and actual values.
+In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together (abdi2007kendall, ).
+Kendall’s tau $\tau$ ranges from $-1$ to $1$, with $\tau=1$ indicating perfect agreement, $\tau=-1$ indicating perfect disagreement, and $\tau=0$ indicating no agreement.
+The Kendall’s tau calculation is computed as follows:
+(2) $\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$
+where $c$ and $d$ denote the numbers of concordant and discordant pairs, $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
+We have separately implemented the two methods in software and found that the PCC computation is much faster than the Kendall’s tau calculation.
+The main reason is that the latter must compute a large full permutation.
+Luckily, when we combined the two methods together, the efficiency and reliability both were enhanced.
+More details will be shown in the following section.
 3. Proposed Improved Two-Step Attack
-It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We will detail the basis of the attack analysis and our proposed three improvements in the following.
-3.1. Basis of Attack Analysis
-The goal of this paper is to explore the physical security of Kyber KEM under random-ciphertext CPA attack.
-The purpose of the attack is to obtain a long-term secret key.
-We assume that an attacker has access to a device which is running Kyber decryption and can enter arbitrary ciphertext into the device.
-In addition, they can also capture the electromagnetic radiation of the device to obtain the power traces.
-In subsection 2.3, we have introduced NTT used in Kyber briefly, which only possesses $n_{\text{th}}$ roots of unity because the modulus polynomial $(x^{n}+1)$ can only be factored into $n/2$ linear polynomials.
-Consequently, multiplication involves the multiplication of linear polynomials rather than simple point-wise multiplication in the NTT domain.
-Due to the properties of incomplete NTT in Kyber, the full secret key of Kyber can be divided into $2k(k=2,3,4)$ groups.
-The test vector leakage assessment (TVLA) methodology can be used to find side-channel leakage points and the t-test method is one of the commonly used methods in TVLA (wu2022efficient, ).
-Several t-test results from previous works have indicated the potential leakage points during the decryption process in Kyber (yang2023chosen, ; cryptoeprint:2022/058, ).
+It should be remarked that the proposed improved two-step attack is illustrated in Fig.
+1.
+We will detail the basis of the attack analysis and our proposed three improvements in the following.
+We assume that an attacker has access to a device which is running Kyber decryption and can only observe ciphertexts entered into the device.
+In addition, they can also capture the power traces of the device.
+Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber (shen2023find, ; yang2023chosen, ).
 One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$.
-Additionally, this operation takes place in the quotient ring $\mathbb{Z}_{q}[x]/(x^{2}-\xi)$.
 In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512.
 Moreover, each calculation of $basemul$ is independent from the others.
-From Listing 1, we can observe that there are two steps related to the partial secret key coefficient $s_{0}$ and three steps related to the partial secret key coefficient $s_{1}$.
-The inputs of a $basemul$ are $s_{0}$, $s_{1}$, $u_{0}$, and $u_{1}$, and outputs are $r_{0}$ and $r_{1}$.
-Five multiplications and two additions are involved in the listing.
-Obviously, the result is related to two secret key coefficients.
+From Listing 1, it is clear that the result depends on two secret key coefficients, $s_{0}$ and $s_{1}$.
 All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber.
-In a practical CPA attack, suppose that the attacker aims to recover the coefficient $s_{0}$, $r_{0}$ can be selected as the intermediate value.
-Similarly, if the attacker aims to recover the coefficient $s_{1}$, $r_{1}$ can be selected as the intermediate value.
-In addition, in order to recover the full secret key coefficients of a group, the attacker would need to perform 128 CPA attacks.
-Therefore, $128\times 2k$ CPA attacks are required for the full secret key.
-Listing 2: Modified HW in the CPA Attack.
-1def HW(r0):
-2 if r0 < 0:
-3 r0 = calculate_complement(r0, 16)
-4 H_W = r0.count("1")
-5 else:
-6 H_W = bin(r0).count("1")
-7 return H_W
-8
-9def calculate_complement(r0, 16):
+In a practical attack, suppose that the attacker aims to recover the coefficient $s_{0}$.
+$r_{0}$ can be selected as the intermediate value.
 3.1. Modified Leakage Model for CPA
-Modeling the power leakage is the basis for a CPA attack, and its accuracy directly determines the success rate of the attack.
-HW model counts the number of “1” in intermediate values, which is one of the most representative linear power models.
-Similarly, HD model records the amount of ”1$\rightarrow$”0” or ”0$\rightarrow$”1” transitions as power leakage.
-The difference of HW and HD model is given as:
-(6) $HD\left(r_{0},s_{1}\right)=HW\left(r_{0}\oplus s_{1}\right).$
-Typically, when simulating the energy consumption of a microprocessor, the HW model is commonly employed (zhao2023side, ), as is the case in this particular study.
-The arithmetic logic units (ALUs) and multiplier units (MU) in ARM processors are often optimized for two’s complement operations.
-Most arithmetic and logical instructions in the processor’s instruction set are designed based on two’s complement calculations.
-Therefore, we fine-tune the HW power leakage model according to this property to directly avoid the complementary false positive.
-We compute the HW of the complement of the intermediate instead of directly computing the HW of the intermediate, as shown in Listing 2, where the function calculate_complement computes the $16bit$ complement.
-The modified power leakage model can map to the hypothetical power consumption values more accurately, thus reducing the number of power traces required for CPA attacks.
-As show in Fig.
-2, correct secret key coefficient in the original HW leakage model is drowned in all the guessed key coefficients, while the coefficient can be picked out in the modified model effectively by using a small number of 15 power traces.
-Figure 2. When $D=15$ during a $basemul$ call, PCC values for all guessed keys of the original HW model (top) and the modified HW model (bottom).
+The Hamming weight (HW) model counts the number of ”1” in intermediate values and is a common linear power model.
+ARM processors’ ALUs are optimized for two’s complement operations, so we adjust the HW (called modified Hamming weight, MHW) to account for this by using the HW of the 16-bit complement of the intermediate values.
+This modification improves the accuracy of the HW model and reduces the number of power traces (denoted as $D$) needed for CPA.
+As shown inside the blue dotted box in Fig.
+2 (a) and (b), the correct key coefficient is obscured using the original model (a) but can be effectively identified with just 15 power traces using the modified model (b).
 3.2. Optional Kendall’s tau for CPA
-The goal of this improvement is to find the correct key coefficients using 15 or even fewer power traces while ensuring the efficiency and success rate of the attack.
-In scenarios where the number of power traces is limited, the calculated PCC value across all guessed keys tends to be exceptionally high.
-Consequently, it becomes challenging to accurately recover even a portion of the secret key coefficients using solely the traditional CPA attack, and attempting to reconstruct 128 or the entire set of secret key coefficients becomes even more daunting.
-Our solution is as follows.
-After computing the PCC correlation matrix for each $basemul$, we set the threshold to 0.9 or even higher, adjusting it based on the values of $\mathbf{P}$.
-For each $basemul$, if there are more than $f$ correlations exceeding the threshold, we consider them as candidate coefficients $\mathbf{k_{h}}=(k_{h_{0}},...,k_{h_{f-1})}$.
-We then map the hypothetical intermediate values of the $f$ candidate coefficients to candidate hypothetical power consumption value matrix $\mathbf{H^{\prime}}_{D\times f}$.
-Next, we calculate the Kendall’s tau between the hypothetical power value matrix $\mathbf{H^{\prime}}$ and the actual power matrix $\mathbf{T}$.
-For each column of $\mathbf{H^{\prime}}$, we compute the Kendall’s tau with each column of the actual power matrix.
-Finally, we obtain a $f\times T$ Kendall coefficient matrix $\mathbf{K_{d}}$, from which we select the candidate key with the highest correlation coefficient as the correct key coefficient for that $basemul$.
-On the other hand, we skip the cases that have higher PCC correlations and smaller candidates.
-PCC is used in the traditional CPA attack to calculate only the linear correlation between the actual power consumption and the hypothetical power consumption, which is not sufficient.
-With only a few power traces, it is possible to directly select some of the correct key coefficients even if only Kendall’s tau is calculated, but the calculation time of Kendall’s tau is 12×12$\times$12 × longer than that of PCC.
-Therefore we combine Kendall’s tau with PCC.
-After calculating PCC, the guessed keys coefficient exceeding a certain threshold is selected as the coefficients of candidate keys.
-Then, the Kendall’s tau values of these coefficients of candidate keys are calculated so that the coefficient with the higher value is accepted.
-In this way, false positives can be quickly eliminated.
-3.4. Trail-and-Error Lattice Attack after CPA
-As mentioned earlier, each group of the secret key consists of 128 secret key coefficients including 128 even-index coefficients or 128 odd-index coefficients.
-As introduced in (kuo2023lattice, ), suppose that an attacker wants to exploit the side-channel leakage point $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$ and has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered.
+When the number of power traces is limited, traditional CPA may face challenges due to high PCC values across guessed keys.
+To address this, we set a PCC threshold (e.g., 0.9) and extract candidate key coefficients with correlations exceeding this threshold.
+We then map these candidates to a hypothetical power consumption matrix and compute Kendall’s tau between this matrix and the actual power traces.
+The candidate with the highest Kendall’s tau value is selected as the correct key coefficient, which helps to reduce false positives.
+This approach combines PCC for initial filtering with Kendall’s tau for more precise selection, despite Kendall’s tau being more computationally demanding.
+Figure 2.
+(a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively.
+(c) shows the ||||PCC|||| values for all guessed keys, while the blue curve in (d) illustrates the occurrence of false positives.
+3.3. Trial-and-Error Lattice Attack after CPA
+As introduced in (kuo2023lattice, ), suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered.
 Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients.
 INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients.
-Similarly, $\mathbf{N}_{B}=[\mathbf{n}_{b_{0}},\ldots,\mathbf{n}_{{b}_{127-sr}}]$, and $\hat{\mathbf{s}}_{iB}=[\hat{\mathbf{s}}_{b_{0}},\ldots,\hat{\mathbf{s}}_{{b}_{127-sr}}]^{\top}$.
+Similarly, $\mathbf{N}_{B}$ and $\hat{\mathbf{s}}_{iB}$ are obtained in the same manner.
 In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known.
@@
 This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller.
-We firstly view it as a bounded distance decoding (BDD)/unique shortest vector problem (uSVP) lattice problem, and specific algorithm to solve the updated LWE problem is given by Algorithm 2 (kannan1987minkowski, ), where the matrixs $\mathbf{BDD}$ and $\mathbf{B}_{kan}$ are summarized as follows:
-(7) $\mathbf{BDD}=\left[\begin{array}[]{cc}\mathbf{I}_{sr}&\mathbf{A}^{\prime}\\ \mathbf{0}&q\mathbf{I}_{n-sr}\\ \end{array}\right],$
+We use Kannan’s embedding technique to solve this problem (kannan1987minkowski, ), where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
 […]
 $\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose.
-This process yields the shortest vector containing the error vector $\mathbf{s}_{i}^{\top}$, denoted as $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$.
-Finally, We obtain the actual secret key with a length of 128 as shown in Step 3 of Algorithm 2.
-But it is worth noting that the norm of vector $\mathbf{w}$ is $\sqrt{\|s_{i}\|^{2}+1}\approx\sqrt{n}\sigma_{s}$, which must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the uSVP problem to be solved.
-For Kyber512, when $sr\geq 39$, it is possible to correctly recover the length of 128 odd/even index coefficients through lattice attack.
+A vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ of length $129$ is obtained by reduction algorithm, which contains the secret key $\mathbf{s}_{i}^{\top}$ of length 128.
+But it is worth noting that the norm of vector $\mathbf{w}$ must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the unique shortest vector problem to be solved.
+For Kyber512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients.
 For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
-In other words, the lattice attack can tolerate at most $89/90$ recovered incorrect secret key coefficients for Kyber512 and Kyber768/1024 respectively, where $38/39$ of the coefficient are randomly selected from $128$ recovery coefficients.
-We elaborate on the theoretical principles of the lattice attack, which provides a fault-tolerant for recovering the full set of 128 keys.
+We elaborate on the theoretical foundation of lattice attacks, which can tolerate error coefficients of recovered key and still successfully retrieve all 128 key.
 Leveraging this method, we propose a more flexible and versatile approach for its application.
-The proposed trail-and-error lattice attack is shown in Listing 3, where $n$ represents the number of CPA attacks, and $\mathbb{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trails, respectively.
-Listing 3: Trail-and-Error Lattice Attack.
+The proposed trial-and-error lattice attack is shown in Listing 2, where $n$ represents the number of CPA, and $\mathbf{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trials, respectively.
+One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA instead of $128$ CPA.
+Certainly, the more CPA are carried out, the higher the success rate of this method will be, but its efficiency will also slow down.
+The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trial-and-error method.
+It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trial and lead to the overall recovery inefficient.
+For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively.
+Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
+Listing 2: Trail-and-Error Lattice Attack.
 1index = [1, 2, ..., 37, 38]
@@
 25def updata_sk():
-Algorithm 2 Kannan’s embedding technique.
-1: Input: An LWE instance $\mathbf{t}=\mathbf{A}\mathbf{s}+\mathbf{s}_{i}\bmod q$
-2: Output: the short error vector $\mathbf{s}_{i}\in\mathcal{R}_{q}^{128}$
-3:Construct the lattice $\Lambda(\text{BDD})$ generated by matrix $\mathbf{BDD}$;
-4:Reduce BDD to the SVP problem by rescaling $\mathbf{BDD}$ to the basis matrix $\mathbf{B}_{kan}$;
-5:Use lattice algorithm (LLL or BKZ) to derive the short vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ from $\mathbf{B}_{kan}$;
-6:return $\mathbf{s}_{i}$
-One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA attacks instead of $128$ CPA attacks.
-Certainly, the more CPA attacks are carried out, the higher the success rate of this method will be, but its efficiency will also slow down.
-The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trail-and-error method.
-It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trail and lead to the overall recovery inefficient.
-For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively.
-Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
-Figure 3. The experimental equipment.
 4. Experiments
-4.1. Experimental Setup
-Our attacks are equally successful for different security levels of Kyber because they use the same $basemul$, whose algorithm is provided by Listing 1.
-For simplicity, we only take Kyber512 into consideration in the following.
-We run the reference implementation of Kyber512 from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz.
-We capture the power traces using a Pico 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe.
-The sampling rate is set to $500$ MSa/s.
-The experimental equipment is shown in Fig.
-3.
-According to the introduction in Section 2.4 that the selection of the intermediate value of CPA attack should be related to the secret key, combined with the analysis mentioned in Section 3.1, we choose $fqmul(k,u_{1})$ as the intermediate value of our attack, and all the intermediate values in the following experiments are the same values.
-Figure 4. When $D=15$, the PCC results during a certain $basemul$ call.
-4.2. Experiment Results
-Step1: CPA Attack
-As aforementioned, the CPA attack is divided into two stages: the capturing stage and the modeling computation stage.
+The proposed attack is applied to Kyber512, and it is noteworthy that our method is also applicable to other security levels due to their use of the same $basemul$.
+We run the reference implementation from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz.
+Power traces were captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a sampling rate set to $500$ MSa/s.
+Based on the analysis in Section 3, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
+Figure 3.
+(a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$.
+Similarly, (c) and (d) show the PCC and Ken results for the same $basemul$ call when $D=11$.
+Step1: CPA
+As aforementioned, the CPA is divided into two stages: the capturing stage and the modeling computation stage.
 Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
-For the capturing stage, as introduced in Section 3.1, we carry out the CPA attack at the NTT operation which calls $basemul$ functions 128 times.
-In our CPA experiments, we input random ciphertexts and capture tremendous electromagnetic radiations during those calls.
+For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions for 128 times.
+In our CPA experiments, we capture random ciphertexts and tremendous data of electromagnetic radiations during those calls.
 To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments.
-The first one is shown in Fig.
-2 of Section 3.2 and we can see that the new model can reinforce the PCC value of the correct key.
-The second one shown in the following is to demonstrate its superiority in identifying the complementary false positive.
-Figure 5. When $D=15$, the PCC results (a) and Ken results (b) in a certain $basemul$. When $D=11$, the PCC results (c) and Ken results (d) in the same $basemul$.
-When the number of random ciphertexts $D$ is set to 15, the relationship of the maximum absolute values $\left|PCC\right|$ to the guess keys is shown in the above of Fig.
-4.
-We can directly pick out the correct key 2280 and its complementary false positive 1049 (equal to $3329-2280$).
-This is a general phenomenon.
-As mentioned in (kuo2023lattice, ), the attacker would run into some problems because the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ both would have high $PCC$ values since $HW(s_{0})$ and $HW(q-s_{0})$ are highly correlated.
-So, we have modified the leakage model as described in Section 3.2 to escape such false positive.
-The relationship of the PCC values to the trace points is shown in the bottom of of Fig.
-4, where there are 3329 curves and the yellow and red curves present whose guess keys are 2280 and 1049, respectively.
-It can be seen that the maximum absolute value of PCC of the complementary false positive has a negative sign.
-We can only focus on the positive PCCs to avoid such false positive.
-We briefly explain the reason of this phenomenon in the following.
-Let $s_{0}$ be the correct one.
-Then $s_{0}^{\prime}=3329-s_{0}$ represents the false positive.
-The $fqmul$ function in $basemul$ is an operation that multiplies two numbers and then is reduced by $3329$ in Kyber.
-The intermediate values for the two guessed key coefficients are summarized as follows:
+The former is shown in Fig.
+2 (a) and (b) of Section 3.1 and the modified model can reinforce the PCC value of the correct key.
+The latter shown in Fig.
+2 (c) and (d) is to demonstrate its superiority in identifying the complementary false positive.
+When the number of random ciphertexts $D$ is set to $15$, the relationship of the maximum ||||PCC|||| to the guess keys is shown in Fig.
+2 (c).
+We can directly pick out the correct key $2280$ and its complementary false positive $1049$ (equal to $3329-2280$), which is a general phenomenon.
+In addition, as the number of power traces decreases, the attacker would run into some problems because both the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ would have high and close PCC values, as the HW$(s_{0})$ and HW$(q-s_{0})$ are highly correlated.
+To escape such false positive, we have modified the leakage model as described in Section 3.1.
+The correspondence between the PCC values of the correct key coefficient and the false positive one is shown in Fig.
+2 (d).
+It is observed that the positive peak value corresponds to the correct key coefficient and the negative peak value of PCC corresponds to the false positive one.
+Consequently, to avoid the false positive one is selected, we should focus solely on the positive peak value of PCC.
+We briefly explain this phenomenon as follows: Let $s_{0}$ be the correct one, and $s_{0}^{\prime}=3329-s_{0}$ be the false positive one.
+The intermediate values for the two guessed key coefficients are summarized below:
 (4) $r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$
@@
 $=-(s_{0}\times u_{1})\,mod\,q.$
-Hypothetical power consumption values $MHW(r_{1})$ and $MHW(r_{1}^{\prime})$ exhibit a negative correlation, while correct $MHW(r_{1})$ will naturally exhibit a positive correlation.
-Obviously, we can figure them out by taking their signs into consideration.
+MHW$(r_{1})$ and MHW$(r_{1}^{\prime})$ show a negative correlation, where the correct MHW$(r_{1})$ is expected to be a positive PCC. Thus, considering their signs allows us to distinguish them.
 To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process.
-Figure 5 (a) shows the points of the maximum absolute PCC results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates $f$ after PCC is equal to 4.
+Fig.
+3 (a) shows the points of the maximum ||||PCC|||| results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates after PCC is equal to 4.
 It can be seen that the correlations of four coefficients $564$, $1687$, $2010$, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to $0.930$ and their correlation values are very close.
-Very close correlations can lead to an inability to distinguish the correct coefficient.
-This is an another kind of false positives.
-If the attacker just picks one of these four at random, the accuracy is only about 25%, which is still regarded as a failure.
+The phenomenon can lead to an inability to distinguish the correct coefficient.
+This is another kind of false positives.
+If the attacker just picks one of the four candidates at random, the accuracy is only about $25$%, which is still regarded as a failure.
 Therefore, we use the method mentioned in Section 3.2 to compensate for such case.
 First of all, we compute the Kendall coefficient matrix of these four candidate coefficients.
-It should be noted that the time for this computation is negligible when comparing to the complete PCC computing.
-Figure 5 (b) shows the points of the four maximum absolute values $\left|Kendall’s\,tau\right|$ to the numbers of power traces.
+It should be noted that the time of this computation is negligible when compared to the complete PCC computing.
+Fig.
+3 (b) shows the values of the four maximum ||||Kendall’s tau|||| when $D=15$.
 It shows that Kendall’s tau expands the differences between correct coefficient and false positives.
-As the number of power traces increases, the value of correct coefficient tends to be more stable while those of false positives decrease quickly.
-We can easily distinguish the correct key 564 when the number of power traces equals 15 and output it as the final result.
-In another word, an attacker can take the correct coefficient 564 for this $basemul$ call using only 15 power traces.
+As the number of power traces increases, the value of correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly.
+We can easily distinguish the correct key $564$ when $D=15$ and output it as the final result.
+In another word, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces.
 Similarly, we reduce the number of power traces to $11$ step by step and implement the PCC and Kendall’s tau computations.
@@
 We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections.
-Note that if we continue reducing the number of power traces, we cannot figure it out any more.
-Step2: Trail-and-Error Lattice Attack
-We conduct the trail-and-error lattice attack on a sixteen-thread server.
+Note that if we continue reducing the number of power traces, we may not be able to figure it out any more.
+Figure 4. The success rate and efficiency of our attack in twenty experiments.
+Step2: Trial-and-Error Lattice Attack
+We conduct the trial-and-error lattice attack using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz.
 The parameter of the $block$ size in the BKZ reduction is set to $50$ and the value of $max\_loops$ to $8$.
-We collect the CPAs of the first sixty $basemuls$ calls.
-Figure 6 shows the relationships of the success rate and time to the number of power traces representing in the blue curve and groups of histograms, respectively.
+We collect the CPA results of the first sixty $basemuls$ calls.
+Fig.
+4 shows the relationships between the success rate/time and the number of power traces represented in the blue curve and groups of histograms, respectively.
 The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times.
 It can be seen that when the number of power traces is no less than $15$, all the experimental attacks are successful.
-Those cases with traces number smaller than 15 can be compensated to some degree by running more trails of the lattice attack.
+Those cases with traces number smaller than $15$ can be compensated to some degree by running more trials of the lattice attack.
 Meanwhile, according to the decomposed time histograms in Fig.
-6, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA attack is almost unchanged although the traces number is doubled.
+4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the traces number is doubled.
 The main reason could be the different utilization ratios of the multi-threading server.
-Note that the CPA computations are independently executed with the different power traces while the trail computations of the lattice attack are executed in serial.
-Therefore, we can summarize that when the success rate is large enough (close to 1), we can only cost 15 power traces for a guess/call and about 9 minutes to recover the full key on Kyber512.
-So, the total number of required power traces is $15\times 60=900$, much smaller than that of the state-of-the-art random ciphertext CPA (yang2023chosen, ) which requires $20\times 128=2560$ power traces.
-If we slightly increasing the traces number of the guess, the required time can be reduced close to 5 minutes, much faster than the recorded 20 minutes reported in (kuo2023lattice, ).
+Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack are executed in serial.
+Therefore, we can summarize that when the success rate is large enough (close/equal to $1$), we can only cost at most $15$ power traces for a guess/call and about $9$ minutes to recover the full key on Kyber512.
+Table 1. SCAs on Kyber to Recover a 128-Value Group of $\mathbf{s}$
+Work Attack Type No Cip Restriction Attack Target Traces Num for a Call Num of Calls
+(xu2021magnifying, ) Pro SPA $\times$ INTT, ME 8$\sim$960 128
+(mu2022voltage, ) Pro TA $\times$ MR 11 128
+(sim2022chosen, ) Non-Pro SPA $\times$ MR 6$\sim$12 128
+(yang2023chosen, ) Non-Pro CPA $\times$ CM 20$\sim$30 128
+(kuo2023lattice, ) Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
+Ours Non-Pro CPA $\checkmark$ CM 11$\sim$15 60
+  • Cip: Ciphertext, Num: Number, Pro: Profile, ME: Message Encoding,
+    MR: Modular Reduction, CM: Coefficient Multiplication.
+Table 1 shows the comparison results of SCAs on Kyber to recover a 128-value group of $\mathbf{s}$.
+The attack types include the profiled SPA (xu2021magnifying, ), profiled TA (mu2022voltage, ), non-profiled SPA (sim2022chosen, ), and non-profiled CPA (kuo2023lattice, ; yang2023chosen, ).
+Our method belongs to the non-profiled CPA, which does not cost extra training step.
+Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme (ravi2024side, ).
+The restriction ciphertext CPA requires oracle-based assistance to be applied to the KEM scheme.
+For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work (yang2023chosen, ) which requires at least $20\times 128=2560$ power traces.
+If we slightly increase the traces number of the guess, the required time can be reduced to approximate $5$ minutes, much faster than the $20$ minutes recorded in (kuo2023lattice, ).
 5. Conclusion
-In this paper, we propose an efficient two-step attack including an enhanced CPA attack and a trail-and-error lattice attack for Kyber.
-In the CPA step, we modify the power leakage model to be more suitable for ARM Cortex-M4 architecture and further filter the candidate keys from the PCC results by using the Kendall’s rank correlation coefficient.
-The accuracy of finding the correct key is significantly improved.
-In the lattice step, we construct a trail-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time.
+In this brief, we propose an efficient two-step attack including an enhanced CPA and a trial-and-error lattice attack for Kyber.
+In traditional CPA, false positives are the main factor affecting coefficients of secret key extraction in experiments (yang2023chosen, )and (kuo2023lattice, ).
+However, by employing the comprehensive CPA step we proposed, we can cleverly avoid false positives while reducing the number of required power traces in the practical attack.
+In the lattice step, we construct a trial-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time.
 Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads.
@@
 Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.
-Figure 6. The success rate and efficiency of our attack in twenty experiments.
-Acknowledgements.
-This work was supported in part by the National Natural Science Foundation of China under Grant 62104097, in part by the Key Research Plan of Jiangsu Province of China under Grant BE2022098, and in part by the Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001.

### v3

@@
-An Improved Two-Step Attack on Kyber
-Abstract.
-After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024.
+An Improved Two-Step Attack on Lattice-based Cryptography: A Case Study of Kyber
+Abstract
+After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by NIST, CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024.
 It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase.
-In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time.
-In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces.
-The CPA is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy.
-In the second step, we adopt the lattice attack to recover s based on the results of CPA.
+In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer power traces and less time.
+In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of power traces.
+The CPA is enhanced by utilizing both Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy.
+In the second step, we adopt the lattice attack to recover $\mathbf{s}$ based on the results of CPA.
 The success rate is largely built up by constructing a trial-and-error method.
-We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine.
-Additionally, in that case, we only cost at most $60$ CPA guess values for a group and 15 power traces for a guess.
-Lattice-based cryptography, CRYSTALS-Kyber, Side-channel attack, Power analysis
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 2024 ACM/IEEE International Conference on Computer-Aided Design; October 27–31, 2024; NEW JERSEY, USAisbn: 978-1-4503-XXXX-X/18/06
-1. Introduction
-WITH the fast development of quantum computing, modern public-key cryptographic algorithms are encountering significant threats.
-Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms.
-By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation (alagic2022status, ).
-CRYSTALS-Kyber is that only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024 (Announcing2024, ).
-It should be marked that the mathematical security of Kyber has been widely recognized by the cryptography community.
-However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security (nagata2021physical, ; zhao2023side, ; ravi2024side, ).
-Kocher et al.
-first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ).
-Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs, especially to the correlation power analysis (CPA).
-Recent years, several CPA attacks on Kyber (ravi2020generic, ; ueno2022curse, ; shen2023find, ; yang2023chosen, ; kuo2023lattice, ) have been proposed by researchers.
-In 2020, Ravi et al (ravi2020generic, ).
-used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time.
-Then, Ueno et al (ueno2022curse, ).
-proposed the side channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key.
-The work in (shen2023find, ) deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber512 at the cost of about 1619 power traces, but there are still 40% error coefficients.
-Contrast to previous works, in (yang2023chosen, ), Yang et al.
-tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces.
-Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber512.
+We deploy the reference implementations of Kyber-512, -768, and -1024 on an ARM Cortex-M4 target board and successfully recover $\mathbf{s}$ in approximately $9\sim 10$ minutes with at most 15 power traces, using a Xeon Gold 6342-equipped machine for the attack.
+Index Terms:
+Lattice-based cryptography, CRYSTALS-Kyber, Side-channel attack, Power analysis.
+I Introduction
+WITH the fast development of quantum computing, modern cryptographic algorithms are encountering significant threats.
+Recognizing this problem, NIST started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms.
+By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation [1].
+CRYSTALS-Kyber is the only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024.
+The mathematical security of Kyber is widely recognized, but its physical security demands urgent attention in the deployment phase [2, 3].
+In recent years, several side-channel attacks (SCAs) on Kyber [4, 5, 6, 7, 8] have been proposed by researchers.
+In 2020, Ravi et al.
+[4] used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber-512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time.
+Then for correlation power analysis (CPA), Ueno et al.
+[5] proposed the side-channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key.
+The work in [6] deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber-512 at the cost of about 1619 power traces, but there are still 40% error coefficients.
+Contrast to previous works, in [7], Yang et al.
+tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber-512, successfully recovering two coefficients of secret key using $20$ or more power traces.
+Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber-512.
 So, they at least need 5120 power traces for a key.
-In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack.
-They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of M-LWE problem.
+In [8], Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack.
+They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of the M-LWE problem.
 The correctness was verified by using computer simulations.
-In this paper, we take the idea of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
-Figure 1. The flowchart of the two-step attack.
+In this brief, we take the idea of [8] as the starting point and develop an improved two-step attack method on Kyber to further validate it in practice and improve its efficiency.
 The proposed two-step attack is summarized in Fig.
@@
 Additionally, we introduce an optional correlation coefficient, the Kendall rank correlation, to address the nonlinearity and robustness issues of the original CPA in real-world attacks.
-In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key, and optimize the parameters of the algorithm.
+In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key and optimize the algorithm’s parameters.
 By employing this two-step attack, we achieve significant improvements in both efficiency and accuracy in practice.
-2. Preliminaries
-2.1. NTT in CRYSTALS-Kyber
+In addition, the attack methods we proposed for the Point-Wise Multiplication (PWM) stage apply to several other lattice-based PQC.
+Dilithium employs the single-point PWM and is more vulnerable to similar attacks.
+The PWM of Falcon works with floating-point numbers, but significant side-channel leakage remains due to secret key and message multiplication [9].
+II Preliminaries
+II-A Kyber and Number Theoretic Transform in Kyber
+Kyber consists of three stages: key generation, encryption, and decryption.
+In the key generation stage, the public key $pk$ is constructed as $\mathbf{A}\mathbf{s}+\mathbf{e}$.
+During the encryption stage, the message $m$ is encrypted into ciphertext $c$.
+In the decryption stage, the receiver utilizes $\mathbf{s}$ to perform corresponding operations to recover the message $m$ from $c$.
+The main operations of decryption are as follows:
+$m\leftarrow Compress_{q}(v-INTT(\boldsymbol{\hat{s}}^{T}\circ NTT(\boldsymbol{u})),1).$ (1)
 Computing polynomial multiplication in NTT domain can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus significantly improving the speed of the entire encryption and decryption process in Kyber.
@@
 Therefore, 256 coefficients are divided into two groups, namely the odd and even groups, for NTT calculations.
-The even coefficients in NTT domian are described as follows:
-(1) $\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$
-where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}$ is a $128\times 128$ matrix with reduced integers.
-Similarly, INTT can be represented in the same way and the inverse matrix is denoted as $N$.
-Listing 1: Base multiplication in Kyber.
-1void basemul(int16_t r[2],
-2 int16_t zeta,
-3 const int16_t s[2],
-4 const int16_t u[2])
-5{
-6 r[0] = fqmul(s[1], u[1]);
-7 r[0] = fqmul(r[0], zeta);
-8 r[0] += fqmul(s[0], u[0]);
-9 r[1] = fqmul(s[0], u[1]);
-10 r[1] += fqmul(s[1], u[0]);
-11}
-2.2. Correlation-Based Attack Techniques
-The conventional CPA attack is a non-invasive side-channel attack that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys.
+The even coefficients in NTT domain are described as follows:
+$\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$ (2)
+where $br$($i$) represents bit-reversal of $i$ and $\mathbf{M}$ is an $128\times 128$ matrix with reduced integers.
+Similarly, the inverse matrix of INTT can be represented as $\mathbf{N}$ in the same way.
+Figure 1: The architecture of the proposed two-step attack.
+II-B Correlation-Based Attack Techniques
+The conventional CPA attack is a non-invasive SCA that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys.
 It identifies the correct key by finding the highest correlation between measured and hypothetical values.
 Pearson correlation coefficient (PCC) is efficient and well-suited for detecting linear relationships between the hypothetical and actual values.
-In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together (abdi2007kendall, ).
+In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together [10].
 Kendall’s tau $\tau$ ranges from $-1$ to $1$, with $\tau=1$ indicating perfect agreement, $\tau=-1$ indicating perfect disagreement, and $\tau=0$ indicating no agreement.
 The Kendall’s tau calculation is computed as follows:
-(2) $\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$
+$\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$ (3)
 where $c$ and $d$ denote the numbers of concordant and discordant pairs, $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
@@
 More details will be shown in the following section.
-3. Proposed Improved Two-Step Attack
+III Proposed Improved Two-Step Attack
 It should be remarked that the proposed improved two-step attack is illustrated in Fig.
 1.
-We will detail the basis of the attack analysis and our proposed three improvements in the following.
 We assume that an attacker has access to a device which is running Kyber decryption and can only observe ciphertexts entered into the device.
 In addition, they can also capture the power traces of the device.
-Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber (shen2023find, ; yang2023chosen, ).
-One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$.
-In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512.
+Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber [6, 7].
+One critical side-channel leakage occurs during the PWM of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$.
+In the clean implementation of pqm4 [11], polynomial multiplication employs 256 $basemuls$ in Kyber-512.
 Moreover, each calculation of $basemul$ is independent from the others.
@@
 In a practical attack, suppose that the attacker aims to recover the coefficient $s_{0}$.
-$r_{0}$ can be selected as the intermediate value.
-3.1. Modified Leakage Model for CPA
-The Hamming weight (HW) model counts the number of ”1” in intermediate values and is a common linear power model.
+$r_{1}$ can be selected as the intermediate value.
+III-A Modified Leakage Model for CPA
+The Hamming weight (HW) model counts the number of “1” in intermediate values and is a common linear power model.
 ARM processors’ ALUs are optimized for two’s complement operations, so we adjust the HW (called modified Hamming weight, MHW) to account for this by using the HW of the 16-bit complement of the intermediate values.
@@
 2 (a) and (b), the correct key coefficient is obscured using the original model (a) but can be effectively identified with just 15 power traces using the modified model (b).
-3.2. Optional Kendall’s tau for CPA
+III-B Optional Kendall’s tau for CPA
 When the number of power traces is limited, traditional CPA may face challenges due to high PCC values across guessed keys.
-To address this, we set a PCC threshold (e.g., 0.9) and extract candidate key coefficients with correlations exceeding this threshold.
+To address this, we set the ||||PCC|||| threshold (e.g., 0.9) based on repeated experimental analysis, with 0.9 performing well in most $basemul$ calls.
 We then map these candidates to a hypothetical power consumption matrix and compute Kendall’s tau between this matrix and the actual power traces.
@@
 This approach combines PCC for initial filtering with Kendall’s tau for more precise selection, despite Kendall’s tau being more computationally demanding.
-Figure 2.
-(a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively.
+Figure 2: (a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively.
 (c) shows the ||||PCC|||| values for all guessed keys, while the blue curve in (d) illustrates the occurrence of false positives.
-3.3. Trial-and-Error Lattice Attack after CPA
-As introduced in (kuo2023lattice, ), suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered.
+III-C Trial-and-Error Lattice Attack after CPA
+As introduced in [8], suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered.
 Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients.
-INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients.
+INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ [8], where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients.
 Similarly, $\mathbf{N}_{B}$ and $\hat{\mathbf{s}}_{iB}$ are obtained in the same manner.
@@
 This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller.
-We use Kannan’s embedding technique to solve this problem (kannan1987minkowski, ), where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
+We use Kannan’s embedding technique to solve this problem [12], where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
 […]
@@
 A vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ of length $129$ is obtained by reduction algorithm, which contains the secret key $\mathbf{s}_{i}^{\top}$ of length 128.
-But it is worth noting that the norm of vector $\mathbf{w}$ must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the unique shortest vector problem to be solved.
-For Kyber512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients.
-For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
+For Kyber-512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients.
+For Kyber-768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
 We elaborate on the theoretical foundation of lattice attacks, which can tolerate error coefficients of recovered key and still successfully retrieve all 128 key.
@@
 The proposed trial-and-error lattice attack is shown in Listing 2, where $n$ represents the number of CPA, and $\mathbf{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trials, respectively.
-One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA instead of $128$ CPA.
-Certainly, the more CPA are carried out, the higher the success rate of this method will be, but its efficiency will also slow down.
-The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trial-and-error method.
-It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trial and lead to the overall recovery inefficient.
-For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively.
-Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
-Listing 2: Trail-and-Error Lattice Attack.
-1index = [1, 2, ..., 37, 38]
-2other_index = [39, 40, ..., n-2, n-1]
-3guess_sk = [sk_bu0, ..., sk_bu38]
-4other_sk = [sk_bu39, ..., sk_bu(n-1)]
-5ken = [k_bu0, ..., k_bu38]
-6other_ken = [k_bu39, ..., k_bu(n-1)]
-7
-8result = Lattice_attack(index, guess_sk)
-9result = (M @ result)
-10if guess_sk == [result[j] for j in index]:
-11 return result
-12else:
-13 for i in count:
-14 updata_index(index)
-15 updata_sk(guess_sk)
-16 result = Lattice_attack(index, guess_sk)
-17 result = (M @ result)
-18 result_target = [result[j] for j in index]
-19 if guess_sk == result_target:
-20 return result
-21 break
-22
-23def Lattice_attack(a, b):
-24def updata_index():
-25def updata_sk():
-4. Experiments
-The proposed attack is applied to Kyber512, and it is noteworthy that our method is also applicable to other security levels due to their use of the same $basemul$.
-We run the reference implementation from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz.
-Power traces were captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a sampling rate set to $500$ MSa/s.
-Based on the analysis in Section 3, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
-Figure 3.
-(a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$.
+In our method, we select 60 basemul calls for the CPA attack and use 38/39 of them as inputs for the lattice attack.
+If the attack fails, the smallest input is replaced with another CPA result.
+This process continues until the attack succeeds or all 60 results are used, representing a trade-off between computational complexity and success rate.
+After extensive experimentation, we found that using 60 results provides similar success to using 128.
+It is also important to note that as the dimension of the lattice problem increases, the computational complexity grows.
+For Kyber, 38/39 results are sufficient to handle the 128 coefficients of a key group.
+However, for Dilithium with a group size of 256, both the lattice dimension and the required number of CPA results will need to increase.
+IV Experiments
+We run the reference implementations of Kyber-512, -768, and -1024 from the pqm4 library [11] with -o compilation optimization on a STM32F407-DISCOVERY target board at $48$ MHz.
+The power traces are captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a $500$ MSa/s sampling rate.
+Based on the analysis in Section III, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
+Figure 3: (a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$.
 Similarly, (c) and (d) show the PCC and Ken results for the same $basemul$ call when $D=11$.
@@
 As aforementioned, the CPA is divided into two stages: the capturing stage and the modeling computation stage.
-Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
-For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions for 128 times.
+Our optimizations are mainly focused on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
+For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions 128 times.
 In our CPA experiments, we capture random ciphertexts and tremendous data of electromagnetic radiations during those calls.
@@
 The former is shown in Fig.
-2 (a) and (b) of Section 3.1 and the modified model can reinforce the PCC value of the correct key.
+2 (a) and (b) of Section III-A and the modified model can reinforce the PCC value of the correct key.
 The latter shown in Fig.
@@
 In addition, as the number of power traces decreases, the attacker would run into some problems because both the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ would have high and close PCC values, as the HW$(s_{0})$ and HW$(q-s_{0})$ are highly correlated.
-To escape such false positive, we have modified the leakage model as described in Section 3.1.
+To escape such false positive, we have modified the leakage model as described in Section III-A.
 The correspondence between the PCC values of the correct key coefficient and the false positive one is shown in Fig.
@@
 It is observed that the positive peak value corresponds to the correct key coefficient and the negative peak value of PCC corresponds to the false positive one.
-Consequently, to avoid the false positive one is selected, we should focus solely on the positive peak value of PCC.
+Consequently, to avoid the false positive being selected, we should focus solely on the positive peak value of PCC.
 We briefly explain this phenomenon as follows: Let $s_{0}$ be the correct one, and $s_{0}^{\prime}=3329-s_{0}$ be the false positive one.
 The intermediate values for the two guessed key coefficients are summarized below:
-(4) $r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$
-(5) $r_{1}^{\prime}=\text{fqmul}(s_{0}^{\prime},u_{1})$ $=((3329-s_{0})\times u_{1})\,mod\,q$
+$r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$ (5)
+$r_{1}^{\prime}=\text{fqmul}(s_{0}^{\prime},u_{1})$ $=((3329-s_{0})\times u_{1})\,mod\,q$ (6)
 $=-(s_{0}\times u_{1})\,mod\,q.$
 MHW$(r_{1})$ and MHW$(r_{1}^{\prime})$ show a negative correlation, where the correct MHW$(r_{1})$ is expected to be a positive PCC. Thus, considering their signs allows us to distinguish them.
+Figure 4: The success rate and efficiency of our attack in twenty experiments.
 To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process.
@@
 It can be seen that the correlations of four coefficients $564$, $1687$, $2010$, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to $0.930$ and their correlation values are very close.
-The phenomenon can lead to an inability to distinguish the correct coefficient.
-This is another kind of false positives.
+The phenomenon can lead to an inability to distinguish the correct one.
+This is another kind of false positive.
 If the attacker just picks one of the four candidates at random, the accuracy is only about $25$%, which is still regarded as a failure.
-Therefore, we use the method mentioned in Section 3.2 to compensate for such case.
+Therefore, we use the method mentioned in Section III-B to compensate for such a case.
 First of all, we compute the Kendall coefficient matrix of these four candidate coefficients.
@@
 3 (b) shows the values of the four maximum ||||Kendall’s tau|||| when $D=15$.
-It shows that Kendall’s tau expands the differences between correct coefficient and false positives.
-As the number of power traces increases, the value of correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly.
+It shows that Kendall’s tau expands the differences between the correct coefficient and false positives.
+As the number of power traces increases, the value of the correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly.
 We can easily distinguish the correct key $564$ when $D=15$ and output it as the final result.
-In another word, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces.
+In other words, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces.
 Similarly, we reduce the number of power traces to $11$ step by step and implement the PCC and Kendall’s tau computations.
@@
 Note that if we continue reducing the number of power traces, we may not be able to figure it out anymore.
-Figure 4. The success rate and efficiency of our attack in twenty experiments.
 Step2: Trial-and-Error Lattice Attack
-We conduct the trial-and-error lattice attack using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz.
+We conduct the trial-and-error lattice attack for a 128-value group of $\mathbf{s}$ (hereinafter referred to as SK-128) using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz.
 The parameter of the $block$ size in the BKZ reduction is set to $50$ and the value of $max\_loops$ to $8$.
@@
 It can be seen that when the number of power traces is no less than $15$, all the experimental attacks are successful.
-Those cases with traces number smaller than $15$ can be compensated to some degree by running more trials of the lattice attack.
+Those cases with trace numbers smaller than $15$ can be compensated to some degree by running more trials of the lattice attack.
 Meanwhile, according to the decomposed time histograms in Fig.
-4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the traces number is doubled.
+4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the trace number is doubled.
 The main reason could be the different utilization ratios of the multi-threading server.
-Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack are executed in serial.
-Therefore, we can summarize that when the success rate is large enough (close/equal to $1$), we can only cost at most $15$ power traces for a guess/call and about $9$ minutes to recover the full key on Kyber512.
-Table 1. SCAs on Kyber to Recover a 128-Value Group of $\mathbf{s}$
+Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack of SK-128 are executed in serial.
+Table I presents the time of implementing the proposed attacks on Kyber at different security levels, using 15 power traces across twenty experiments.
+Kyber-512, -768, and -1024 have 4, 6, and 8 SK-128s, respectively, with each SK-128 being independent of the others.
+This allows us to recover the different SK-128s in parallel.
+The attack method we propose can recover the full secret key on different security levels of Kyber, with the success rate remaining consistent.
+The reason for the slight increase in time taken at the high security level is the high parallelism of Step1 and the duration of Step2 fluctuates around 5 minutes.
+It is worth noting that reducing the parallelism will increase the time required for Kyber-768 and Kyber-1024, but the success rate remains unaffected.
+TABLE I: Time Consumed to the Proposed Attacks on Kyber by Using 15 Power Traces in Twenty Experiments ($n=60$)
+Security Level Step1 Step2 Total Time ($min$)
+Time / Thread Time / Thread
+SK-128 4.76 / $16\times 1$ 4.67 / $1\times 1$ 9.43
+Kyber-512 5.01 / $16\times 4$ 4.66 / $1\times 4$ 9.67
+Kyber-768 5.14 / $16\times 6$ 5.00 / $1\times 6$ 10.14
+Kyber-1024 5.30 / $16\times 8$ 5.33 / $1\times 8$ 10.63
+TABLE II: SCAs on Kyber to Recover SK-128
 Work Attack Type No Cip Restriction Attack Target Traces Num for a Call Num of Calls
-(xu2021magnifying, ) Pro SPA $\times$ INTT, ME 8$\sim$960 128
-(mu2022voltage, ) Pro TA $\times$ MR 11 128
-(sim2022chosen, ) Non-Pro SPA $\times$ MR 6$\sim$12 128
-(yang2023chosen, ) Non-Pro CPA $\times$ CM 20$\sim$30 128
-(kuo2023lattice, ) Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
+[13] Pro SPA $\times$ INTT, ME 8$\sim$960 128
+[14] Pro TA $\times$ MR 11 128
+[15] Non-Pro SPA $\times$ MR 6$\sim$12 128
+[7] Non-Pro CPA $\times$ CM 20$\sim$30 128
+[8] Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
 Ours Non-Pro CPA $\checkmark$ CM 11$\sim$15 60
@@
     MR: Modular Reduction, CM: Coefficient Multiplication.
-Table 1 shows the comparison results of SCAs on Kyber to recover a 128-value group of $\mathbf{s}$.
-The attack types include the profiled SPA (xu2021magnifying, ), profiled TA (mu2022voltage, ), non-profiled SPA (sim2022chosen, ), and non-profiled CPA (kuo2023lattice, ; yang2023chosen, ).
+Table II shows the comparison results of SCAs on Kyber to recover SK-128.
+The attack types include the profiled SPA [13], profiled TA [14], non-profiled SPA [15], and non-profiled CPA [7, 8].
 Our method belongs to the non-profiled CPA, which does not cost extra training step.
-Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme (ravi2024side, ).
+Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme [2].
 The restriction ciphertext CPA requires oracle-based assistance to be applied to the KEM scheme.
-For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work (yang2023chosen, ) which requires at least $20\times 128=2560$ power traces.
-If we slightly increase the traces number of the guess, the required time can be reduced to approximate $5$ minutes, much faster than the $20$ minutes recorded in (kuo2023lattice, ).
-5. Conclusion
-In this brief, we propose an efficient two-step attack including an enhanced CPA and a trial-and-error lattice attack for Kyber.
-In traditional CPA, false positives are the main factor affecting coefficients of secret key extraction in experiments (yang2023chosen, )and (kuo2023lattice, ).
-However, by employing the comprehensive CPA step we proposed, we can cleverly avoid false positives while reducing the number of required power traces in the practical attack.
-In the lattice step, we construct a trial-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time.
-Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads.
-Because the $basemul$ function called by different security levels of Kyber is the same, our work can be directly applied to the other parameters.
+For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work [7] which requires at least $20\times 128=2560$ power traces.
+If we slightly increase the traces number of the guess, the required time can be reduced to approximately $5$ minutes, much faster than the $20$ minutes recorded in [8].
+V Conclusion
+In this brief, we propose an efficient two-step attack for Kyber, combining an enhanced CPA and a trial-and-error lattice attack.
+Experimental results show that the proposed attack can accurately recover the full secret key of Kyber-512, -768, and -1024 in about $9\sim 10$ minutes with about 15 power traces.
 Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.
+It is worth noting that our proposed attack method can be applied to implementations on hardware like FPGAs by slightly modifying the leakage model.
+To mitigate these attacks, algorithm-level protections such as shuffling [16] and masking [17] can be employed to protect the secret key, though these methods may involve trade-offs in efficiency and effectiveness.